OpenStudy (anonymous):

Domain and range of f(x)=square root of (49-x^2)

OpenStudy (jdoe0001):

\(\bf \sqrt{49-x^2}\) what values for "x" will make the radical quantity 0?

OpenStudy (anonymous):

7 and -7. I got confused cause i am not sure whether to look at it as a circle or as a square root function

OpenStudy (jdoe0001):

7 and -7 are correct, so that's your domain, "x" will have to be between -7 and 7 if it becomes bigger than that, you'd end up with a negative inside the radical and that'd make it an imaginary number

OpenStudy (anonymous):

whatever is inside of the square root has to be > or = to 0. So \[49-x ^{2}\ge 0\] Resolve to get x alone and you'll have the domain. As for the range, y will always be > or = to what? hint, look at the parent function.

OpenStudy (jdoe0001):

as far the Range, "y", well the radical when "x" becomes 7 or -7 will turn to 0 and when "x" is in between -7 and +7, 49 will be the bigger value and thus positive how far up positive? well, if we set "x = 0", you'd be left with \(\bf \sqrt{49-0}\) which yields no more than 7 or -7